OpenStudy (anonymous):

Identify, graph, and state the symmetries for each polar equation. Write the scale that you are using for the polar axis. 1) r=9cos(5theta) 2) r=2costheta

OpenStudy (anonymous):

I honestly don't really know how to do this stuff

OpenStudy (perl):

hint cos(-Θ)= cos(Θ)

OpenStudy (perl):

If the equation is unchanged when you replace r with -r you have symmetry with respect to origin (with respect to pole). If the equation is unchanged when you replace Θ with -Θ you have symmetry with respect to x axis (with respect to polar axis). If the equation is unchanged when you replace r with -r and Θ with -Θ you have symmetry with respect to y axis (with respect to Θ = π/2).

OpenStudy (anonymous):

@perl, I just don't get what it's asking me to do.

OpenStudy (anonymous):

@Haseeb96 can you help?

OpenStudy (anonymous):

@amistre64 help?

OpenStudy (anonymous):

@roz_mar23 do you understand this?

OpenStudy (dan815):

hi :)

OpenStudy (dan815):

illl help uuuu

OpenStudy (anonymous):

hello @dan815 !!

OpenStudy (anonymous):

thank you!

OpenStudy (dan815):

give me one min i have to wash my plate

OpenStudy (anonymous):

for sure

OpenStudy (dan815):

okkkk

OpenStudy (dan815):

lets start with the base equation \[r=cos(\theta )\]

OpenStudy (dan815):

you know how cos(theta)=x/r?

OpenStudy (anonymous):

no. I don't know anything about precalc. this is my last packet for school and I'm learning it online and it doesn't make sense.

OpenStudy (dan815):

ok sure, dont worry it just comes from the circle, do you remember how cos theta = adj/hyp from the Soh Cah Toa stuff

OpenStudy (anonymous):

yes that much i think i understand

OpenStudy (dan815):

okay well take a look at the x y coordinate system

OpenStudy (dan815):



OpenStudy (dan815):

first of all this is the relationship between r and theta with x and y

OpenStudy (dan815):